OpenStudy (ranya99):

Free Fall Quiz-assume air resistance is ignored 1. Luke drops a pile of roof shingles from the top of a roof located 8.52m above the ground. Determine the time required for the shingles to reach the ground. 2. Rex throws his mother’s crystal vase vertically upwards with an initial velocity of 26.2m/s. Determine the height to which the vase will rise above the initial height. At what velocity is the vase moving when Rex catches it. 3. An object falls from the Transco Tower in Houston and takes 15seconds to reach the ground. What is the velocity at impact and how tall is the building?

OpenStudy (anonymous):

Ima go one at a time beecausse i may have to go

OpenStudy (ranya99):

alright

OpenStudy (ranya99):

continuing the question 4. An orange is tossed upward at 21m/s. What is the velocity of the orange 3.5s later? What is the height of the orange at this time? Is the orange still travelling up or is it travelling down? 5. A tennis ball is tossed up off a building with a velocity of 22m/s. It takes 6.4s to reach the ground. How high is the building and what is the maximum height of the tennis ball?

OpenStudy (anonymous):

so the distance the object falls is given by the equation 1/2gt^2 which means that 8.52=1/2(9.8xt^2) we can change the subject to solve for time; \[\sqrt{\frac{ 2s }{ g }}=t\] and now we substitute the values into the equation to get, t=1.3seconds

OpenStudy (anonymous):

okay so we know the rate of deceleration of the vase that is assumeing rex lives on earth that is -9.8 m/s and we also know the initial velocity which is 26.2m/s so to find the time it takes to reach the peak height we just need to devide 26.2 by 9.8 to get 2.67seconds. to find the max hieghtwe multiply 26.2 into the time and we get 70m

OpenStudy (anonymous):

okay so we know the rate of deceleration of the vase that is assumeing rex lives on earth that is -9.8 m/s and we also know the initial velocity which is 26.2m/s so to find the time it takes to reach the peak height we just need to devide 26.2 by 9.8 to get 2.67seconds. to find the max hieghtwe multiply 26.2 into the time and we get 70m

OpenStudy (anonymous):

now the rest of the questions are basicallythe same two concepts so you can answering them by substituteing the values into the equations i gave you.

OpenStudy (anonymous):

oh, and thank you for 'fanning me' :P

OpenStudy (ranya99):

alright, thank you so much :) & no problem